OpenStudy (anonymous):

The circumference of a circle is 5 inches. What's the radius?

OpenStudy (campbell_st):

radius is 5/2pi

OpenStudy (anonymous):

Ummm...ok. How do I do that?

OpenStudy (campbell_st):

C = 2pi x r then r = C/2pi if C = 5 then r = 5/2pi...

OpenStudy (anonymous):

Ok.. Thank you.